# Key Words.: cosmology: distance scale – gravitational lensing: strong – Galaxy: kinematics and dynamics – Galaxies: elliptical and lenticular, cD1 Introduction

### v1

@@ 5.3 Verifying polynomial orders setting
For a sample of galaxies or multiple spatial bins within a galaxy, identified by the subscript , one would have to compute the average across the sample , then compute the weighted average of it across templates index using the global BIC,
and the systematic uncertainty on the mean
For a small number of stellar libraries, as in the case discussed here, we recommend applying Bessel’s correction:
In the end, it is important to verify that the combination of polynomial orders is around the minimum within the uncertainty. In Figures 15 and 16, we show as an example the dependency of the inferred stellar velocity dispersion as a function of polynomial orders for the MUSE sample. The left-hand side panel shows the BIC-weighted average of the three clean stellar libraries, while the right-hand side panel shows the straight average. In all cases, the results change much less than 1% for polynomial orders around our fiducial values. Bessel-corrected estimators have been used.
5.4 Summary of Recipe
Here, we summarize our suggested recipe step-by-step to obtain accurate and precise stellar velocity dispersion –
@@ 5.4 Summary of Recipe
Use data with sufficient signal-to-noise ratio. If the signal-to-noise ratio is too low, the inference will not only be more noisy, but can also be biased. The precise threshold depends on the wavelength coverage and spectral resolution of the dataset and also on the target accuracy and precision. This can be estimated by simulations or checked post-facto using this recipe itself, varying the threshold in SNR.
@@
Use clean stellar template libraries. Popular stellar libraries include spectra that may be interesting for other applications, but have insufficient quality for precise and accurate determination. We provide our “clean” list for MILES, XSL, and Indo-Us through GitHub, with the caveat that additional cuts/masking may be required depending on the wavelength and quality of the data under consideration.
@@
Run initial tests on additional and multiplicative polynomial orders to identify the range in which the results are stable. Depending on the wavelength range and quality of the data, it may be necessary to marginalize over polynomial orders while computing the total systematic error budget.
@@
(Optional) If the systematic error associated with the template library is smaller than 0.5%, and a better estimate of the systematic error is needed, compute the additional term by evaluating residual scatter by marginalizing over polynomial orders and small changes in the wavelength range.
@@ 6 Results
In this section, we summarize and discuss our main results, namely the precision and accuracy of the stellar velocity dispersion measurements obtained by applying our recipe to our datasets. We stress that, whereas the recipe is general and can be applied to any dataset, the precision and accuracy that can be achieved are specific to each dataset, and depend on the signal-to-noise ratio, resolution, wavelength range, and intrinsic properties of the galaxy spectra under consideration. We will, therefore, first discuss each dataset separately and then draw some general conclusions. The key numbers are summarized in Table 2. For convenience, we report both the values obtained with our optimal recipe and also those obtained with a straight average of the three clean stellar libraries.
@@ 6.1 MUSE
The BIC-weighted average systematic error is much smaller than our target of 1%, even with Bessel correction. This is, of course, only the error associated with template choice. Other potential sources of error at the level of 0.2-0.5% are the wavelength range, instrumental resolution, and polynomial order, and those should be investigated if one wants to push accuracy to the level of 0.3-0.4%.
Not surprisingly, the velocity dispersion has positive covariance across the sample: templates that yield higher (lower) velocity dispersion do so consistently for all galaxies. Importantly, the amplitude of the covariance is also well below 1%.
If one were to ignore the information from the BIC and weigh the three clean stellar libraries equally, the uncertainties associated with the stellar template would increase to the values indicated in the lower part of Table 2. The increase is noticeable and strengthens the importance of adopting our proposed recipe. However, even with this suboptimal weight, the systematic errors and off-diagonal elements of the covariance matrix remain sub-percent.
@@
SDSS | 3.64% | 1.45% | 1.30% | 1.78% | 1.60% | 0.333 | 0.333 | 0.333 |
6.1.1 KCWI
The results for KCWI are very similar to those obtained for MUSE, with the main difference being that Indo-US is now more strongly favored with respect to both XSL and MILES. This may be due to the superior SNR of the data, reflected in the sub-percent statistical errors. Since the data tell us that they clearly prefer Indo-US, then the residual systematics obtained with our procedure are vanishingly small, and Bessel correction is crucial to obtain a realistic estimate of the systematic error. For comparison, a straight average of the three clean stellar libraries leaves systematic residual uncertainties and covariance below 1%.
@@ 6.3 SDSS
The results for SDSS are qualitatively in line with the other instruments, even though the accuracy and precision is lower, as expected given that the SNR of the data is significantly lower. As expected, average random errors are higher than MUSE and KCWI, between 3-4%. Remarkably, even with the lower data quality, systematic errors and covariance are just above 1% for BIC weighting and still below 2% with the straight average. The most significant difference is that the SDSS spectra prefer XSL over Indo-US, whereas the opposite is true for MUSE and KCWI. MILES is always disfavored. Our suggested explanation for this difference is that the longer and redder wavelength coverage of the SDSS spectra compared with MUSE and KCWI (3600-5350 Å vs 3600-4500 Å). For example, the Indo-US flux calibration may be better in the blue but less accurate than XSL in the longer and redder wavelength range. The stellar populations that dominate the redder absorption features could also be better described by the XSL library. In contrast, those that dominate in the blue could be better described by Indo-US. This difference highlights the importance of running all the clean stellar libraries.
@@ 6.4 General results
We now summarize results that are general and valid across the samples considered here. We also expect them to be valid for other samples with data of comparable quality. First, once properly cleaned stellar libraries are considered, residual systematics on stellar velocity dispersion due to the template choices are sub-percent. Second, the uncertainty can be further reduced by taking advantage of the information content of the data to select which set of templates better describes the data, according to the procedure described in Section 5. We note that the exact outcome of this process will depend on the data quality and quantity (that set the statistical significance) but also on the wavelength range covered by the data and by the stellar populations in the target galaxies (that may be best described by a specific library or not). Third, the covariance between different galaxies or different spatial bins is also at the sub-percent level. Fourth, there is no single clean stellar library that is preferred by all datasets, so it is important to use all three.
@@ 7 Summary
We carried out a detailed study of systematic uncertainties associated with stellar velocity dispersion measurements in samples of massive elliptical galaxies at redshift . We use spectra with SNR30-160 Å-1 obtained with three different state-of-the-art instruments, VLT-MUSE, Keck-KCWI, and JWST-NIRSpec. For comparison with previous works, we also study a sample of relatively high SNR spectra from SDSS (SNR15 Å-1). We consider three state-of-the-art empirical stellar libraries (Indo-US, MILES, and XSL) as templates. For completeness, we also investigate synthetic libraries that are enhanced (sMILES) and composite stellar population synthesis (SSP) models that have been used in previous studies.
Our main results can be summarized as follows:
@@
It is very important to use “clean” stellar libraries, i.e., removing spectra that have defects, such as gaps, poor correction of slit losses, atmospheric absorption and differential diffraction, etc. Not properly cleaned stellar libraries can result in differences of order a few percent and sometimes up to 10-15%.
@@
SSP and synthetic -enhanced libraries can produce results that differ by a few percent with respect to clean stellar libraries, and we discourage their use for high-precision stellar kinematics. For the SSP libraries, we interpret the bias as being due to incomplete coverage of parameter space. This yields insufficient flexibility to reproduce the detailed star formation and chemical enrichment history of massive elliptical galaxies to our desired level of precision. For the -enhanced libraries, the bias is likely due to known limitations in the modeling of stellar atmospheres (Knowles et al. 2023).
@@
Once clean stellar libraries are considered, the residual uncertainty related to the choice of templates is sub-percent for all the primary datasets considered here (MUSE/KCWI/NIRSpec). Even for the lower SNR SDSS spectra they are below 2%. The off-diagonal elements of the covariance matrix between galaxies in a sample and between spatial bins within a galaxy are positive and also sub-percent.
@@
Other sources of residual uncertainty related to the orders of the polynomials used to model the continuum are shown to be of order 0.2-0.5% for our dataset.
In conclusion, we have shown that adopting the method suggested here leads to a substantial improvement in the precision and accuracy of stellar kinematics measurements, as well as the estimation of the residual uncertainties. While the final precision and accuracy will depend on the properties of the dataset, including wavelength range, the methodology is fully general and can be applied to any dataset. Adopting our procedure enables the standardization and optimization of the measurement of stellar velocity dispersions. This is a crucial step if one wants to use them for cosmography, and will be adopted as standard by the TDCOSMO collaboration.
To facilitate the adoption of our method by third parties, the list of templates and the scripts used in this analysis will be made publicly available after the acceptance of this manuscript through the GitHub repository available at URL https://github.com/TDCOSMO/KINEMATICS_METHODS.
Looking ahead, it is important to note that – When using a measured to estimate velocity moments for dynamical models – there are additional sources of potential systematics and model-dependent effects that must be evaluated, mitigated, or corrected, to ensure that accuracy on translates to our target of on H0. Such considerations include how accurately a Gaussian line-of-sight velocity distribution (LOSVD) approximates the true velocity moment and the corresponding accuracy and precision in mass estimates derived from such models, as well as triaxiality and projection effects (Huang et al. 2025, in preparation). These will be addressed in future papers, building upon the key result of this paper, i.e. the ability to accurately and precisely determine the observable quantity .

### v2

@@ 5.3 Verifying polynomial orders setting
For a sample of galaxies or multiple spatial bins within a galaxy, identified by the subscript , we would have to compute the average across the sample , then compute the weighted average of it across templates index using the global BIC,
and the systematic uncertainty on the mean,
For a small number of stellar libraries, as in the case discussed here, we recommend applying Bessel’s correction,
In the end, it is important to verify that the combination of polynomial orders is around the minimum within the uncertainty. In Figures 12 and 13, we show as an example the dependency of the inferred stellar velocity dispersion as a function of polynomial orders for the MUSE sample. The left-hand side panel shows the BIC-weighted average of the three clean stellar libraries, while the right-hand side panel shows the straight average. In all cases, the results change much less than 1% for polynomial orders around our fiducial values. Bessel-corrected estimators have been used.
5.4 Summary of the recipe
We summarize our suggested recipe step by step to obtain accurate and precise stellar velocity dispersion below.
@@ 5.4 Summary of the recipe
Use data with sufficient S/N. If the S/N is too low, the inference will not only be more noisy, but can also be biased. The precise threshold depends on the wavelength coverage and spectral resolution of the dataset and also on the target accuracy and precision. This can be estimated by simulations or checked post facto using this recipe itself, varying the threshold in S/N.
@@
Use clean stellar template libraries. Popular stellar libraries include spectra that might be interesting for other applications, but have insufficient quality for precise and accurate determination. We provide our “clean” list for MILES, XSL, and Indo-Us through GitHub, with the caveat that additional cuts and masking might be required depending on the wavelength and quality of the data under consideration.
@@
Run initial tests on additional and multiplicative polynomial orders to identify the range in which the results are stable. Depending on the wavelength range and quality of the data, it might be necessary to marginalize over polynomial orders while computing the total systematic error budget.
@@
In this optional step, if the systematic error associated with the template library is smaller than 0.5%, and a better estimate of the systematic error is needed, the additional term is computed by evaluating residual scatter by marginalizing over polynomial orders and small changes in the wavelength range.
@@ 6 Results
In this section, we summarize and discuss our main results, namely, the precision and accuracy of the stellar velocity dispersion measurements achieved by applying our recipe to our datasets. We stress that, while the recipe is general and can be applied to any dataset, the precision and accuracy that can be achieved are specific to each dataset and depend on the S/N, resolution, wavelength range, and intrinsic properties of the galaxy spectra under consideration. Therefore, we first discuss each dataset separately and draw some general conclusions. The key numbers are summarized in Table 2. For convenience, we report both the values obtained with our optimal recipe and also those obtained with a straight average of the three clean stellar libraries.
@@ 6.1 MUSE
The BIC-weighted average systematic error is much smaller than our target of 1%, even with Bessel correction. This is, of course, only the error associated with template choice. Other potential sources of error at the level of 0.2-0.5% are the wavelength range, instrumental resolution, and polynomial order, and those should be investigated if we want to push the accuracy to the level of 0.3-0.4%.
Not surprisingly, the velocity dispersion has positive covariance across the sample: templates that yield higher (lower) velocity dispersion do so consistently for all galaxies. Importantly, the amplitude of the covariance is also well below 1%. If we were to ignore the information from the BIC and weigh the three clean stellar libraries equally, the uncertainties associated with the stellar template would increase to the values indicated in the lower part of Table 2. The increase is noticeable and strengthens the importance of adopting our proposed recipe. However, even with this suboptimal weight, the systematic errors and off-diagonal elements of the covariance matrix remain on the sub-percent level.
@@
| SDSS | 3.30% | 1.30% | 1.15% | 1.60% | 1.40% | 0.333 | 0.333 | 0.333 |
6.2 KCWI
The results for KCWI are very similar to those obtained for MUSE, with the main difference being that Indo-US is now more strongly favored with respect to both XSL and MILES. This might be due to the superior S/N of the data, reflected in the sub-percent statistical errors. Since the data inform us that they clearly prefer Indo-US, then the residual systematics obtained with our procedure are vanishingly small and a Bessel correction is crucial to obtain a realistic estimate of the systematic error. For comparison, a straight average of the three clean stellar libraries leaves systematic residual uncertainties and covariance below 1%.
@@ 6.4 SDSS
The results for SDSS are qualitatively in line with the other instruments, even though the accuracy and precision is lower, as expected given that the S/N of the data is significantly lower. As expected, average random errors are higher than MUSE and KCWI, between 3-4%. Remarkably, even with the lower data quality, systematic errors and covariance are below 1.5% for BIC weighting and still below 2% with the straight average. The most significant difference is that the SDSS spectra prefer XSL over Indo-US, whereas the opposite is true for MUSE and KCWI. MILES is always disfavored. Our suggested explanation for this difference is that the longer and redder wavelength coverage of the SDSS spectra compared with MUSE and KCWI (3600-5350 Å vs 3600-4500 Å). For example, the Indo-US flux calibration might be better in the blue but less accurate than XSL in the longer and redder wavelength range. The stellar populations that dominate the redder absorption features could also be better described by the XSL library. In contrast, those that dominate in the blue could be better described by Indo-US. This difference highlights the importance of running all the clean stellar libraries.
@@ 6.5 General results
We now summarize results that are general and valid across the samples considered here. We also expect them to be valid for other samples with data of comparable quality. First, once properly cleaned stellar libraries are considered, residual systematics on stellar velocity dispersion due to the template choices are sub-percent. Second, the uncertainty can be further reduced by taking advantage of the information content of the data to select which set of templates better describes the data, according to the procedure described in Section 5. We note that the exact outcome of this process will depend on the data quality and quantity (that set the statistical significance) but also on the wavelength range covered by the data and by the stellar populations in the target galaxies (that might be best described by a specific library or not). Third, the covariance between different galaxies or different spatial bins is also at the sub-percent level. Fourth, there is no single clean stellar library that is preferred by all datasets, so it is important to use all three.
We carried out additional tests of our proposed recipe with mock samples constructed from the clean stellar template libraries of Indo-US and XSL, demonstrating that the method is able to accurately recover the true velocity dispersion to sub-percent accuracy. The details of the mock sample and results are included in Appendix B.
@@ 7 Summary
We carried out a detailed study of systematic uncertainties associated with stellar velocity dispersion measurements in samples of massive elliptical galaxies at redshift . We used spectra with S/N30-160 Å-1 obtained with three different state-of-the-art instruments: VLT-MUSE, Keck-KCWI, and JWST-NIRSpec. For comparison with previous works, we also studied a sample of relatively high-S/N spectra from SDSS (S/N15 Å-1). We consider three state-of-the-art empirical stellar libraries (Indo-US, MILES, and XSL) as templates. For completeness, we also investigated synthetic libraries that are enhanced (sMILES) and composite stellar population synthesis (SSP) models that have been used in previous studies. Our main results can be summarized as follows:
@@
It is very important to use “clean” stellar libraries, that is, to remove spectra that have defects, such as gaps, poor correction of slit losses, atmospheric absorption, differential diffraction, and so on. Improperly cleaned stellar libraries can result in differences of order a few percent and sometimes up to 10-15%.
@@
We find that SSP and synthetic -enhanced libraries can produce results that differ by a few percent with respect to clean stellar libraries and we discourage their use for high-precision stellar kinematics. For the SSP libraries, we interpret the bias as being due to incomplete coverage of parameter space. This yields insufficient flexibility to reproduce the detailed star formation and chemical enrichment history of massive elliptical galaxies to our desired level of precision. For the -enhanced libraries, the bias is likely due to known limitations in the modeling of stellar atmospheres (Knowles23).
@@
Once clean stellar libraries are considered, the residual uncertainty related to the choice of templates is sub-percent for all the primary datasets considered here (MUSE/KCWI/NIRSpec). Even for the lower S/N SDSS spectra they are below 2%. The off-diagonal elements of the covariance matrix between galaxies in a sample and between spatial bins within a galaxy are positive and also sub-percent.
@@
Other sources of residual uncertainty related to the orders of the polynomials used to model the continuum are shown to be on the order of 0.2-0.5% for our dataset.
In conclusion, we have shown that adopting the method suggested here leads to a substantial improvement in the precision and accuracy of stellar kinematics measurements, as well as the estimation of the residual uncertainties. While the final precision and accuracy will depend on the properties of the dataset (including the wavelength range), the methodology is fully general and can be applied to any dataset. Our procedure enables the standardization and optimization of the measurement of stellar velocity dispersions. This is a crucial step required for using them for cosmography, and it will be adopted as standard by the TDCOSMO collaboration.
Looking ahead, it is important to note that when using a measured to estimate velocity moments for dynamical models, there are additional sources of potential systematics and model-dependent effects that must be evaluated, mitigated, or corrected, to ensure that a accuracy on translates to our target of on H0. Such considerations include how accurately a Gaussian line-of-sight velocity distribution (LOSVD) approximates the true velocity moment and the corresponding accuracy and precision in mass estimates derived from such models, as well as triaxiality and projection effects tdcosmo_xxi_triaxiality. These aspects will be addressed in future papers, building upon the key result of this paper, namely, the ability to accurately and precisely determine the observable quantity, .
8 Data availability
To facilitate the adoption of our method by third parties, the list of templates and the scripts used in this analysis has been made publicly available through the GitHub repository available at URL https://github.com/TDCOSMO/KINEMATICS_METHODS. We have also developed the squirrel pipeline to implement our method. The pipeline was first introduced in Shajib25 and has been made publicly available at https://github.com/ajshajib/squirrel. squirrel is built on the penalized PiXel Fitting (pPXF) software package (Cappellari17; Cappellari23).